a diagonal of a square is 3 inches longer than a side of a square. find the length of the side
Answer by edjones(4080)

Let x=any side, x+3=the diagonal
a^2+b^2=c^2 Pythagoras
2x^2=(x+3)^2
2x^2=x^2+6x+9
x^2-6x-9=0 subtract the right side from the left.
x=7.243..." length of the side.
.
Ed